OpenStudy (anonymous):

Find the value of x for a to be parallel to b. m<1 = 117 and m<2 = 6x – 12.

OpenStudy (anonymous):

Diagram?

OpenStudy (anonymous):

Angles 1 and 2 must add up to 180

OpenStudy (anonymous):

117 + 6x - 12 = 180 6x + 105 = 180 6x = 75 x = 12.5

OpenStudy (anonymous):

that's what I got but the answers are all in fractions.

OpenStudy (ranga):

Is 25/2 one of the answers?

OpenStudy (anonymous):

Yes, 25/2 = 12.5

OpenStudy (anonymous):

okay thank you! yeah it is but I just didn't know how to change it.

OpenStudy (ranga):

you are welcome.